OpenStudy (annie12m):

Dean drank 1280 mL of water on Monday. On Tuesday he increased his water intake by 82%. How much water did Dean drink on Tuesday?

OpenStudy (annie12m):

@ScarlettFarra2000

OpenStudy (jingle535):

Hello again, to solve a problem like this you need to multiply 1280 by 82% or 0.82 then add it back to the original amout (1280)

OpenStudy (annie12m):

1049.6

OpenStudy (jingle535):

Yes now add that to 1280, because its 82% more

OpenStudy (scarlettfarra2000):

he is drinking additional 82% to his 100%. (100% is 1280 mL) so he is taking 182% of the 1280mL 182% of something , >>> something times 1.82. \[1280mL \times 1.82\] you will have: f(x)=1280(1.82)x, as your function, for any day x, with whater f(x) You find 82% of 1280 Long story short its 2329.6.

OpenStudy (annie12m):

1049.6 + 1280 = 2329.6

OpenStudy (annie12m):

k thanks everyone

OpenStudy (jingle535):

No problem :)

OpenStudy (jingle535):

Sorry scarlet you had to type that all out, ill give you a medal ;)

OpenStudy (scarlettfarra2000):

It's okay as long as annie12m got the answer